OpenStudy (anonymous):

A race horse took 2 2/5 minutes to run 1 1/2 miles. How long would it take the horse to run one mile? A. 9/10 minute B. 1 1/3 minutes C. 1 1/10 minutes D. 1 2/5 minutes

OpenStudy (anonymous):

C. 1 1/10 minutes - Dragonemperor

OpenStudy (study312):

Do you know how to make 2 2/5 improper?

OpenStudy (study312):

since you are offline I'll show you

OpenStudy (study312):

2 • 5 + 2 = 12/5. 1 • 2 + 1 = 3/2. 12 • 3 = 36 2 • 5 = 10, 36/10. 36/10 in simplest form is 18/5.